In my opinion, we ought to stop making our own drums and accept that outside suppliers offer, said Wim Niewindt, managing
Answer #1

1.Financial Advantage of Buying Drums from outside supplier = Relevant cost of producing – cost of buying

= Variable Cost of making + supervision cost + rental cost – cost of buying

= (10.50+7.5*70% + 1.50*70%)*65,000 + 58,500 + 175,500 – 18*65,000

= $156,000

i.e. advantage

52.Financial Advantage = 195,000*16.8 + 58,500 + 175,500 – 18*195,000

= $0

i.e. indifferent

3. Financial Advantage = 16.8*225,000 + 58,500 + 175,500 – 18*225,000

= -$36,000

i.e. disadvantage